OpenStudy (anonymous):

sin^2x/1-cosx=1+cosx Trig identity proofs

OpenStudy (anonymous):

sin^2x = 1-cos^2x

OpenStudy (anonymous):

L =sin^2x / 1-cosx P=1+cosx L= 1-cos^2x/1-cosx = (1-cosx)(1+cosx)/1-cosx = 1+cosx =p L=P

OpenStudy (anonymous):

1-cos^2x is the the difference of squares (1-cosx)(1+cosx)

OpenStudy (anonymous):

1-cosx will cancel leaving 1+cosx=1+cosx